### v1

@@ Abstract
AI is beginning to transform food science across the entire innovation cycle in Figure 3–from ingredient design, formulation development, and fermentation and production to texture analysis, sensory properties, manufacturing, and recipe generation. Yet, the field’s true potential lies not only in optimizing today’s processes, but also in charting a bold trajectory for the future of food as a programmable, predictive science. Building on recent advances, we envision three major areas that define the frontier of AI for sustainable food production:
Future Materials for Food. Treating food as a programmable biomaterial opens new opportunities to design ingredients from first principles. Figure 8 highlights how properties such as texture, functionality, stability, and sensory perception emerge from multiscale interactions among proteins, carbohydrates, lipids, and their processing conditions [20].
Figure 8: Future Materials for Food. Treating food as a programmable multiscale material–from molecular interactions, supramolecular structures and assemblies, and networks to the final food product–opens new opportunities to design ingredients from first principles. In this biomaterials approach, properties such as texture, functionality, stability, and sensory perception emerge from multiscale interactions among proteins, carbohydrates, lipids, and their processing conditions.
@@
We may be able to optimize edible films and gels for controlled nutrient release; or engineer plant-based meat analogues with tailored mechanical and sensory performance to endow such new components with both exceptional sensory experiences and nutritional value. Advances in protein design, including AlphaFold and diffusion-based generative models [102], suggest the possibility of de novo edible proteins with tunable performance. Early examples from cellular agriculture–such as engineered porcine fat cells with modified fatty acid composition [103, 104] or bovine muscle cells enriched in antioxidants or vitamins [105]–demonstrate the promise, while microbial and fungi-based fermentation highlight a complementary path. Mycelium fermentation, for instance, can upcycle agricultural side-streams such as brewers’ spent grain or fruit pomace, converting low-value byproducts into high-quality protein and fiber-rich ingredients with unique textures and link innovation with circularity and sustainability [106]. An opportunity lies ahead in integrating molecular design with tunable processing, including fermentation and upcycling, to create predictive platforms that generate new food materials with tailored properties while reducing waste.
@@
Automation and Self-Driving Labs. Empirical trial-and-error approaches that use rheological measurements, sensory panels, and compositional analysis have traditionally dominated food science [107]. These methods, while valuable, are slow, costly, and difficult to scale. Generative AI, robotics, and high-throughput experimentation point toward self-driving labs that autonomously design, test, and refine food materials. This embeds the design-build-test-learn principle directly into experimentation, enabling systematic exploration across scales and accelerating innovation. Recent advances in automated model discovery with custom-designed constitutive neural networks, show how AI can autonomously identify the best models, parameters, and even experimental protocols directly from data [108]–a method already successfully applied to compare plant-based and animal meats [69], deli products [109], and fungi-based steak [106]. Agentic AI systems such as Sparks [110], BioDiscoveryAgent [111], and the Virtual Lab [112] already demonstrate the capacity to generate hypotheses, perform simulations, and report results without human intervention. In food, such systems could integrate genome-scale metabolic models, digital twins of bioreactors, and in-line quality sensors [113, 67]. A fully automated discovery pipeline that links synthesis, characterization, and modeling has the potential to accelerate innovation in sustainable, personalized, and scalable food production.
@@
Deep Reasoning Models. AI models that can reason will become a key part of self-driving lab infrastructure. Figure 9 highlights our vision for next-generation AI systems for food. We anticipate that the emerging discipline of AI for Food will benefit from ongoing rapid progress in general-purpose reasoning models as well as agentic systems that can autonomously design and perform tasks with the aid of additional tools [114]. While current generative models can propose new molecules or recipes, they remain limited by static datasets and lack the ability to reason about feasibility, manufacturability, or emergent properties. Bridging this gap requires an AI system that can generate new protein sequences, simulate their folding and mechanical behavior, and reason about emergent sensory properties, e.g., flavor, taste - all within a self-improving, interpretable loop. Deep reasoning models represent the next frontier: AI systems that move beyond pattern recognition to actively hypothesize, simulate, and generate new data. These models would integrate protein sequences, structural mechanics, metabolic fluxes, and sensory outcomes, reasoning across scales to predict how molecular composition translates into taste, texture, and stability [115, 20]. Crucially, they could bridge the simulation-to-reality gap by iteratively validating predictions against experimental data and refining models in real time. These approaches could leverage automated model discovery to identify the best experiments and models for alternative protein products [68], seamlessly integrate the results into finite element simulations [116] to simulate real-life chewing processes and aromas from cooking, and predict sensory perception directly from physical principles [69]. By embedding sustainability constraints, cultural diversity, and personalized health data, such systems could design foods that are not only functional but also socially and nutritionally relevant. The ultimate vision is a closed-loop, reasoning-based AI that autonomously proposes, evaluates, and refines edible materials—transforming food science into a predictive, design-driven discipline for future foods.
From Vision to Implementation. To translate the emerging opportunities of AI-enabled food systems into lasting impact, the community must unite around shared goals, standards, and responsibilities. Specifically, we will need to: i) establish common standards and open data repositories for the field with aligned incentives, security, and risk mitigation; ii) overcome the language barrier between wet-lab and computational scientists; iii) develop clear evaluation frameworks to match AI tools to food challenges and measure outputs; and iv) build consumer trust and prepare for evolving regulatory landscapes. But this field can only progress alongside parallel advancements and such a paradigm shift will require support across all sectors: Researchers must push real-time imaging, spectroscopy, and miniaturized hardware forward to generate high-quality, high-throughput data streams. Engineers and scientists need to expand automation and robotics so that self-driving laboratories and closed-loop cycles become practical. Developers of AI must improve compute sustainability to ensure models run energy-efficiently and align with environmental goals. Computational social scientists must work to understand the impact of current sustainable proteins and guide AI-enabled sustainable protein developers toward sensory and nutritional profiles and educational messages likely to increase displacement effects. Policymakers, ethicists, and scientists together must build governance frameworks in ethics, regulation, and cultural inclusivity so that AI-enabled food systems earn consumer trust and deliver equitable benefits across geographies. Funders and institutions must invest in the needed infrastructure–shared repositories, standards, and sensing and automation hardware–to support the interdisciplinary centers and fellowships needed to carry out this work. Together, we must ground this field in responsible innovation by designing evaluation frameworks, building regulatory clarity, and embedding transparency, interpretability, and inclusivity. These infrastructural investments echo the lessons from drug discovery and protein design, where progress accelerated only once community benchmarks, standards, and testbeds were in place. Moving forward, we need to think critically about engaging AI, recognizing the environmental footprint, and ensuring that automation responsibly complements existing scientific endeavors.

### v2

@@ Conclusions
AI is beginning to transform food science across the entire innovation cycle, from ingredient design, formulation development, and fermentation and production to texture analysis, sensory properties, manufacturing, and recipe generation (Fig. 2). Yet, the field’s true potential not only lies in optimizing today’s processes, but also in charting a bold trajectory for the future of food as a programmable, predictive science. Building on recent advances, we envision four major areas that define the frontier of AI for sustainable food production:
@@
Future materials for food innovation. Treating food as a programmable biomaterial creates opportunities to design ingredients from first principles: texture, functionality, stability, and sensory perception emerge from multiscale interactions among proteins, carbohydrates, lipids, and processing conditions [20]. This perspective enables the design of edible films and gels with controlled nutrient release and plant-based meat analogues with tailored mechanical and sensory performance. Advances in protein design, including AlphaFold and diffusion-based generative models [92], point toward de novo edible proteins with tunable functionality. Early work in cellular agriculture, such as engineered fat and muscle cells with modified composition [93], demonstrates this potential, while microbial and fungi-based fermentation offer complementary routes. Mycelium fermentation can upcycle agricultural side-streams such as brewers’ spent grain or fruit pomace into protein- and fiber-rich ingredients with distinctive textures and clear links to circularity [94]. A key opportunity lies in integrating molecular design with tunable processing, including fermentation and upcycling, to build predictive platforms that generate new food materials with targeted properties and reduced waste.
Scientific machine learning for food systems. Scientific Machine Learning (SciML) [95] hardwires physical, chemical, and biological principles directly into AI architectures and provides a foundation for trustworthy and scalable food innovation. Physics-Informed Neural Networks (PINNs) [96] enforce conservation laws in data-sparse settings [97] such as fermentation and extrusion. Constitutive Artificial Neural Networks (CANNs) [98] embed mechanical laws to discover physically interpretable models of food texture and rheology [64]. Sparse Identification of Nonlinear Dynamics (SINDy) [99] discovers compact governing equations from experimental data and can reveal the kinetics of ingredient interactions or microbial growth. The opportunity ahead is to couple physics-based ML with data-driven models to achieve interpretable, generalizable predictions under sparse data and enable mechanism-aware food design.
@@
Automation and self-driving labs. Empirical trial-and-error methods based on rheology, sensory panels, and compositional analysis have long shaped food science, but they remain slow, costly, and hard to scale. Generative AI, robotics, and high-throughput experimentation enable self-driving labs that design, test, and refine food materials within a closed-loop design–build–test–learn cycle, accelerating systematic exploration across scales. Recent advances in automated model discovery with custom-designed constitutive neural networks, show how AI can autonomously identify the best models, parameters, and even experimental protocols directly from data [98]–a method already successfully applied to compare plant-based and animal meats [64], deli products [100], and fungi-based steak [94]. Agentic AI systems such as Sparks [101], BioDiscoveryAgent [102], and the Virtual Lab [103] demonstrate the capacity to generate hypotheses, perform simulations, and report results without human intervention. In food, such systems could integrate genome-scale metabolic models, digital twins of bioreactors, and in-line quality sensors [104, 62]. A fully automated discovery pipeline that links synthesis, characterization, and modeling would substantially accelerate innovation in sustainable, personalized, and scalable food production.
@@
Deep reasoning models. AI models with reasoning capabilities will become central to self-driving lab infrastructure (Fig. 4). The emerging discipline of AI for Food will benefit from rapid progress in reasoning models and agentic systems that design and execute tasks with external tools [105]. Current generative models propose molecules and recipes, but they rely on static datasets and struggle to assess feasibility, manufacturability, or emergent properties. Closing this gap requires systems that generate protein sequences, simulate folding and mechanics, and predict sensory outcomes such as flavor and texture within an interpretable, self-improving loop. Deep reasoning models define the next frontier beyond pattern recognition, and instead hypothesize, simulate, and generate data. They integrate protein sequences, structural mechanics, metabolic fluxes, and sensory responses to predict how molecular composition translates into taste, texture, and stability [20]. They can reduce the simulation-to-reality gap by validating predictions against experiments and updating models in real time. Such approaches can guide experiment selection [63], integrate results into finite element simulations of oral processing [106], and predict sensory perception from physical principles [64]. By embedding sustainability constraints, cultural diversity, and personalized health data, these systems can design foods that meet functional, nutritional, and societal goals. The ultimate goal is a closed-loop, reasoning-driven platform that designs, tests, and refines food materials with predictive accuracy across scales.
Limitations and risks. This review uses sustainable proteins as a testbed for AI-enabled food systems and focuses on post-harvest product development, from ingredients to recipes. It does not cover upstream agriculture, supply chains, packaging, safety testing, or retail analytics, domains that also adopt AI and merit dedicated reviews. Benchmarking remains limited by heterogeneous datasets, proprietary silos, and weak metadata standards, which constrain reproducibility and global representation. AI integration also introduces technical and societal risks, including limited transferability, domain shift, and bias from sparse data. In food systems, these risks carry direct safety implications: inconsistent outputs may yield unsafe formulations, opaque optimization may hide allergen or nutritional trade-offs, and automated processes raise safety and biosecurity concerns. Responsible use of AI requires transparent benchmarks, physics-informed constraints, human-in-the-loop validation, regulatory alignment, and inclusive data design. AI must augment expert oversight to ensure safe, scalable, and sustainable innovation.
Outlook. Taken together, these four areas define a path toward a predictive and design-driven science of food. AI can link molecular composition to structure, processing, and perception and integrate these levels into closed-loop systems that design, test, and refine foods across scales. From programmable biomaterials and physics-informed models to self-driving laboratories and reasoning-based AI, the field now has the foundations to move from empirical optimization to mechanism-aware design. Realizing this vision will require interoperable data standards, rigorous benchmarking, and platforms that integrate experiments, simulations, and learning. Progress will also depend on human-centered design that embeds cultural context, nutrition, and safety into AI systems. If these challenges are addressed, AI will not only accelerate food innovation, but establish a new paradigm in which food design becomes predictive, scalable, and aligned with both human and planetary health.